# FAIR Enough: How Can We Develop and Assess a FAIR-Compliant Dataset for Large Language Model Training?

### 19 Jan 2024 v1

@@ 4.2.1 Data Collection and Curation for LLM Development
In accordance with FAIR data principles, our dataset has been meticulously structured to enhance its findability through the inclusion of comprehensive metadata. This collection, sourced from diverse channels (news feeds, website) between January and May 2023, comprises over 50,000 diverse entries, significantly enhancing its accessibility and discoverability. In our efforts to engage with current social discourse, we utilized a variety of feeds and hashtags. The selected hashtags for data curation include #MediaBias, #SocialJustice, #GenderEquality, #RacialInjustice, #CulturalDiversity, #AgeismAwareness, #ReligiousTolerance, and #EconomicDisparity.
@@
The dataset’s metadata includes information such as the dataset title, description, authors, date of creation, version, and keywords that reflect the dataset’s content and purpose. Additionally, the metadata includes keywords (LLMs, Training, Biases, News Media, NLP) for easy retrieval in these portals and specifies the data type, be it textual, numerical, or otherwise.
The Gunning Fog Index is a readability test designed to assess the level of education required to understand a piece of writing Jacobs (2023). A higher Gunning Fog Index value indicates that the text is more complex and requires a higher level of education to understand it. Figure 5 displays the distribution of Gunning Fog Index scores for our dataset, with a normal distribution centered around a mean score of 7.79. The range of scores suggests that most texts are suitable for readers with an eighth-grade education level. This data helps us evaluate the readability of these texts to match the intended audience’s comprehension abilities.
@@
Figure 5: Histogram of the Gunning Fog Index scores for our dataset. The x-axis denotes the Gunning Fog Index scores, reflecting text complexity, and the y-axis represents the number of samples with each score.
@@
The annotation process of this dataset is a sophisticated blend of human expertise and algorithmic efficiency, involving five domain experts and two ML-trained students. Such an annotation approach is also supported in the recent literature Liu et al. (2023); Shu et al. (2023), as being better approach compared to traditional crowd-sourcing. We employ an active learning algorithm, specifically ‘Uncertainty Sampling’ Nguyen et al. (2022) for bulk annotations. In this approach, the ML model identifies data points where human expertise can impactfully clarify ambiguities. Periodic expert reviews and inter-annotator agreement (IAA) checks, evidenced by a Cohen’s Kappa score above 0.75, ensure high-quality annotations of our data. Few IIA agreement across main bias dimensions is shown in Appendix .3.
@@
For interoperability, the data is formatted for seamless integration for various ML tasks, such as binary and multi-label classifiers, and question answering systems, as detailed in Table 3. This enables researchers to use this data in diverse analytical contexts, facilitating cross-domain research and development. The dataset schema and examples on dataset formats are given in Appendix .2.
@@
The robustness of our dataset for LLM training is established through a dual-stage quality assessment: automated analysis for initial screening of 10,000 statistically significant entries, and expert review of 500 stratified random samples for deeper insights. This method ensures the dataset’s accuracy, completeness, and consistency, crucial for LLM training, as detailed in Table 4.
Table 4: Data Quality Analysis Summary
Aspect Automated Analysis Expert Review
Accuracy 96.5% Contextually accurate
Completeness 93.0% High with occasional updates
Consistency 98.0% Uniform across samples
@@ 4.2.2 Model Training and Algorithm Development
The model training and algorithm development phase is crucial for adhering to FAIR principles, particularly interoperability and reusability. We strategically deploy various LLMs, fine-tuned for tasks ranging from sentiment analysis and question answering to debiasing, and develop them with a modular design for enhanced reusability across projects.
@@
We prepare API documentation that provides details for every aspect of training and development, including parameter settings and algorithm modifications. This transparency facilitates scientific validation and progress Jacobsen et al. (2020b). We also foster community collaboration by sharing our findings, models, and tools. The goal is to contribute to public repositories, and participating in AI research platforms.
@@ 4.2.3 Model Evaluation and Validation
Figure 6: Heatmap of labels across main dimensions
@@
Table 5: Comprehensive Analysis of Language Model Performance and Impact of Debiasing. The first section details performance metrics of various models on different tasks, fine-tuned for optimal settings. The second section highlights the effect of debiasing on toxicity, bias classification, and sentiment scores.
@@
The Table 5 demonstrates that all models initially exhibit strong classification performance. However, after the debiasing process, a noticeable decline in classification performances is observed across these models when applied to the debiased dataset. This suggests that the original training was heavily reliant on biased data, and the introduction of debiased data disrupts the model’s classification ability, leading to a reduction in accuracy. This result aligns very much with the recent safety benchmarking in recent works Touvron et al. (2023), which shows there is trade-off highlights the impact of debiasing on model performance. Overall,these models’ evaluation and validation phase, emphasize transparent reporting and public availability of models and the detailed results, aligns with the FAIR principles of Reusability and Accessibility.
@@ 4.2.4 Deployment and Ongoing Monitoring
In the post-deployment phase, we emphasize the FAIR principles of Accessibility, Reusability, and Findability. Our approach includes providing comprehensive documentation and clear usage guidelines to enhance the model’s findability and accessibility. We employ version control systems to ensure easy tracking and retrieval of the latest model versions, supporting both findability and reusability.
To facilitate user engagement and model effectiveness, we offer extensive training materials and support, making the models more accessible to a diverse user base. Regular reporting of performance metrics offers insights into the models’ ongoing effectiveness and areas for improvement, aiding in their reusability. We also provide integration support with various tools and platforms, ensuring that our models can be easily adopted in different technological contexts.
These practices collectively ensure that our deployed models adhere to the FAIR principles, maintaining their utility and efficacy in a dynamic, user-centric environment.
@@ 4.2.5 Community Engagement and Collaborative Development
In our efforts to foster community engagement and collaborative development, we align with the FAIR principles of Findability, Accessibility, Interoperability, and Reusability. We provide open-source development to facilitate collaborative innovation. Our datasets are provided in accessible formats such as JSON and CSV and are indexed with rich metadata. This approach ensures that our LLMs are developed in environments conducive to collaborative progress.
The datasets are disseminated under the Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license Creative Commons (2023), aligning our work with the FAIR principles of data stewardship.
@@
Table 6: FAIR Compliance in Bias Identification Dataset and LLM Development
FAIR Dimension FAIR Compliance Criteria Met Remarks (Implementation Details)
@@
This approach ensures that our LLMs are technologically advanced, ethically sound, and conducive to a robust, inclusive future in AI development.
@@ 5.1 Limitations of the FAIR Data Principles in Addressing Data Challenges
The FAIR data principles are a crucial step towards more open and efficient use of data in research and beyond. However, their effectiveness is limited by challenges related certain factors discussed below:
  • Data Quality Oversight: Although the FAIR principles make data more accessible and usable, these princicples do not inherently guarantee the quality or validity of the data. There is no built-in mechanism to ensure the data’s accuracy, completeness, or reliability. This lack of quality control can lead to the proliferation of poor-quality data and can impact research and decision-making processes.
  • Ethical and Privacy Concerns: While FAIR principles promote accessibility and sharing, they do not sufficiently address the complex ethical implications and privacy issues related to sensitive data. In areas such as patient health records or personal information, the unrestricted sharing of data could lead to privacy violations and ethical dilemmas.
  • Resource and Cost Implications: Implementing FAIR principles can be resource-intensive. It requires infrastructure, tools, and expertise to make the process happen. This can be particularly challenging for smaller institutions or developing countries with limited resources, potentially widening the digital divide.
  • One-Size-Fits-All Approach: The FAIR principles is not one-size-fits-all solution for all scientific disciplines. Different fields may require tailored approaches to data management that the generalized FAIR guidelines do not fully accommodate.
  • Lack of Universally Accepted Standards: While the FAIR principles provide a framework, there is still a lack of universally accepted standards and metrics to evaluate how ’FAIR’ a dataset is. This ambiguity can lead to inconsistent applications of the principles.
  • Technical Challenges in Interoperability: Data from different sources often exist in various formats and standards, it is difficult to make seamless integration across different systems and platforms. The over-emphasis on data sharing might overshadow other important aspects of data stewardship, such as data privacy.
  • Intellectual Property and Commercial Interests: FAIR principles may conflict with intellectual property rights and commercial interests. Making data freely accessible and reusable might undermine proprietary research or commercial advantages, leading to resistance from certain sectors.Adhering to the FAIR principles must also be balanced with compliance with legal and regulatory frameworks, which can vary significantly across regions.
  • Training and Awareness: Effective implementation of FAIR principles requires a high level of awareness and training among data practitioners. The lack of such training can be a significant barrier to the adoption of these principles.
5.2 Strategies for Mitigating Limitations and Enhancing Data Utility
@@
In addressing the limitations of FAIR Data Principles and enhancing data utility, a multifaceted strategy is essential. This involves refining data quality through rigorous validation processes, ensuring accuracy and reliability. Simultaneously, balancing data accessibility with ethical and privacy considerations is crucial, there is need to employ advanced anonymization techniques and strict privacy protocols. Emphasizing data preservation and sustainability alongside sharing, broadens the scope of stewardship. Integrating legal and regulatory compliance is vital for the robust and ethical utilization of data in research and practice.
@@ 5.3 Limitation of the Study and Future Perspectives
Our Bias-Identification dataset, designed to identify and mitigate known biases, may inadvertently overlook emerging biases, thereby emphasizing the need for ongoing revision and vigilant monitoring May et al. (2019); Raza (2021). In tandem with the rapid development of LLMs, this necessitates adaptable datasets and methodologies to accommodate new architectures and ensure wide-ranging applicability Wilson et al. (2023). Concurrently, scaling the dataset to match the complexity of advanced LLMs, while adhering to ethical standards, emerges as a substantial challenge. This endeavor is compounded by the imperative to mitigate biases in model interpretations, bolster interpretability, and ensure robust data privacy and security Zhao et al. (2023a); Chen et al. (2023). Therefore, future research directions should consider creating dynamic mechanisms for dataset updates, advancing bias mitigation techniques, aligning datasets with evolving LLM technologies, exploring scalable dataset maintenance solutions, and broadening the spectrum of LLM applications. Integral to this journey is the development of comprehensive ethical guidelines and the fostering of collaborative research, both of which are pivotal for the responsible evolution of LLM technology.

### 23 Jan 2024 v2

@@ 4.2.1 Data Collection and Curation for LLM Development
In line with the FAIR data principles, our dataset is meticulously structured to enhance its Findability, featuring comprehensive metadata to facilitate easy discovery. This dataset, sourced from a variety of channels such as news feeds and websites between January and May 2023, encompasses over 50,000 filtered entries. For data curation, we utilized various feeds and hashtags, including #MediaBias, #SocialJustice, #GenderEquality, #RacialInjustice, #CulturalDiversity, #AgeismAwareness, #ReligiousTolerance, and #EconomicDisparity, to ensure a wide representation of social issues. This diversity of news articles not only broadens accessibility but also enriches the dataset relevance to current social discourses.
@@
The detailed metadata includes crucial information such as dataset title, description, authors, date of creation, version, and keywords that reflect the dataset scope and purpose. These keywords (e.g., LLMs, Training, Biases, News Media, NLP) are strategically chosen to enhance the dataset retrieval efficiency across various research portals. Additionally, the metadata specifies the data type, whether textual, numerical, or otherwise, further aligning with the Interoperability principle by facilitating data integration across different systems.
This readability test is crucial for determining the level of education required to comprehend our dataset texts. As illustrated in Figure 5, the distribution of Gunning Fog Index scores in our dataset shows a normal distribution with a mean score of 7.79. This indicates that the majority of our texts are suitable for readers with at least an eighth-grade education level. Such readability analysis is instrumental in aligning our dataset with the FAIR principle of Accessibility, ensuring that the content is comprehensible to the intended audience, thus enhancing the usability of the dataset in training LLMs.
@@
[Uncaptioned image] Figure 5: Histogram of the Gunning Fog Index on FAIR-Complaint Dataset. The x-axis denotes the Gunning Fog Index scores, reflecting text complexity, and the y-axis represents the number of samples with each score.
Table 3: Data Quality Analysis Summary Aspect Automated Analysis Expert Review Accuracy 96.5% Contextually accurate Completeness 93.0% High with occasional updates Consistency 98.0% Uniform across samples
@@
The annotation of this dataset is a sophisticated combination of human expertise and algorithmic efficiency, involving a team of five domain experts and two machine learning-trained students. This approach aligns with recent studies Liu et al. (2023); Shu et al. (2023) that advocate for it as a superior alternative to traditional crowd-sourcing methods. For bulk annotation, we utilized an LLM-based model (specifically, the paid API of GPT-3 turbo), followed by human review to effectively resolve ambiguities. This process is enhanced by periodic expert reviews and inter-annotator agreement (IAA) checks. The quality of our annotations is evidenced by a Cohen’s Kappa score consistently above 0.75. Examples of IAA agreements across main bias dimensions are provided in Appendix .3, illustrating the robustness of our annotation methodology.
The robustness of our dataset for LLM training is established through a dual-stage quality assessment: automated analysis for initial screening of 10,000 statistically significant entries, and expert review of 500 stratified random samples for deeper insights. This method ensures the dataset accuracy, completeness, and consistency, crucial for LLM training, as detailed in Table 3.
@@
For Interoperability, the data is formatted for seamless integration for various ML tasks, such as binary and multi-label classifiers, question answering (QA) system and debiased language generation task (debiasing), as detailed in Table 4. This enables researchers to use this data in diverse analytical contexts, facilitating cross-domain research and development. The dataset schema and examples on dataset formats are given in Appendix .2.
@@ 4.2.2 Model Training and Algorithm Development
The model training and algorithm development phase is crucial for adhering to FAIR principles, particularly interoperability and reusability. We strategically deploy various LLMs, fine-tuned for tasks ranging from sentiment analysis and QA to debiasing (language generation), and develop them with a modular design for enhanced Reusability across projects.
@@
We prepare API documentation that provides details for every aspect of training and development, including parameter settings and algorithm modifications. This transparency facilitates scientific validation and progress Jacobsen et al. (2020). We also foster community collaboration by sharing our findings, models, and tools. The goal is to contribute to public repositories, and participating in AI research platforms.
@@ 4.2.3 Model Evaluation and Validation
Figure 6: Heatmap Visualization: the prevalence and intensity of different types of biases, such as ageism, gender, and political, across various classifications like bias, non-biased, toxic, and sentiment within a dataset.
@@
Table 5: Comprehensive Analysis of Language Model Performance and Impact of Debiasing: the first section details performance metrics of various models on different tasks, fine-tuned for optimal settings. The second section highlights the effect of debiasing on toxicity, bias classification, and sentiment scores.
@@
The results in the table demonstrate that all models initially exhibit strong classification performance. However, after the debiasing process, a noticeable decline in classification performances is observed across these models when applied to the debiased dataset. This suggests that the original training was heavily reliant on biased data, and the introduction of debiased data disrupts the model’s classification ability, leading to a reduction in accuracy. This result aligns very much with the recent safety benchmarking in recent works Touvron et al. (2023), which shows there is a trade-off highlighting the impact of debiasing on model performance.
Overall, these models’ evaluation and validation phase, emphasize transparent reporting and public availability of models and the detailed results, aligns with the FAIR principles of Reusability and Accessibility.
@@ 4.2.4 Deployment and Ongoing Monitoring
In the post-deployment phase, we emphasize the FAIR principles of Accessibility, Reusability, and Findability. Our approach includes providing comprehensive documentation and clear usage guidelines to enhance the model findability and Accessibility. We employ version control systems to ensure easy tracking and retrieval of the latest model versions, supporting both findability and Reusability.
To facilitate user engagement and model effectiveness, we offer extensive training materials and support, making the models more accessible to a diverse user base. Regular reporting of performance metrics offers insights into the models’ ongoing effectiveness and areas for improvement, aiding in their reusability. We also provide integration support with various tools and platforms, ensuring that our models can be easily adopted in different technological contexts. These practices collectively ensure that our deployed models adhere to the FAIR principles, maintaining their utility and efficacy in a dynamic, user-centric environment.
@@ 4.2.5 Community Engagement and Collaborative Development
In our efforts to foster community engagement and collaborative development, we align with the FAIR principles effectively. We provide open-source development to facilitate collaborative innovation. Our datasets are provided in accessible formats such as JSON and CSV and are indexed with rich metadata. This approach ensures that our LLMs are developed in environments conducive to collaborative progress. The datasets are disseminated under the Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license Creative Commons (2023), aligning our work with the FAIR principles of data stewardship.
@@
Table 6: FAIR Compliance in Dataset and LLM Development
Principle FAIR Compliance Criteria Met Remarks (Implementation Details)
@@ 5.1 Limitations of the FAIR Data Principles in Addressing Data Challenges
The FAIR data principles mark a significant step towards enhancing the openness and efficiency of data usage in research and related fields. Yet, their effectiveness is constrained by several factors. Primarily, while the FAIR principles improve data accessibility and usability, they do not automatically guarantee data quality or validity. The absence of mechanisms for ensuring accuracy, completeness, or reliability could lead to the spread of low-quality data, negatively impacting research and decision-making. Additionally, FAIR principles, though encouraging data sharing, may not adequately address the ethical and privacy concerns associated with sensitive data like health records or personal information, which could result in privacy violations.
Furthermore, the implementation of FAIR principles requires considerable resources, including necessary infrastructure and expertise. This poses a significant challenge, especially for smaller institutions or those in developing countries, potentially exacerbating the digital divide. The broad application of FAIR principles may not be suitable for all scientific disciplines, as different fields might require tailored data management approaches not fully encompassed by these general guidelines. The effectiveness of these principles also depends heavily on the awareness and training of data handlers, where a lack of such training can be a major barrier to adoption. Moreover, the absence of universally accepted standards for evaluating ’FAIRness’ leads to an inconsistent application of these principles.
Additionally, the diversity of data formats creates technical challenges in achieving seamless interoperability, where the emphasis on data sharing might overshadow other vital aspects of data stewardship, such as privacy considerations. Finally, the principles can conflict with intellectual property rights and commercial interests. The concept of free data access and reuse might challenge proprietary research, leading to resistance from certain sectors. This necessitates compliance with various legal and regulatory frameworks, adding another layer of complexity to the application of FAIR principles.
Strategies for Mitigating Limitations and Enhancing Data Utility
@@
In addressing the limitations of FAIR principles and enhancing data utility, a multifaceted strategy is essential. This involves refining data quality through rigorous validation processes, ensuring accuracy and reliability. Simultaneously, balancing data accessibility with ethical and privacy considerations is crucial, there is need to employ advanced anonymization techniques and strict privacy protocols. Emphasizing data preservation and sustainability alongside sharing, broadens the scope of stewardship. Integrating legal and regulatory compliance is vital for the robust and ethical utilization of data in research and practice.
@@ 5.2 Limitation of the Study and Future Perspectives
Our Fair-Compliant dataset, designed to identify and mitigate known biases, may inadvertently overlook emerging biases, thereby emphasizing the need for ongoing revision and vigilant monitoring May et al. (2019). Concurrently, scaling the dataset to match the complexity of advanced LLMs, while adhering to ethical standards, emerges as a substantial challenge. This endeavor is compounded by the imperative to mitigate biases in model interpretations, bolster interpretability, and ensure robust data privacy and security Zhao et al. (2023a); Chen et al. (2023). Therefore, future research directions should consider creating dynamic mechanisms for dataset updates Wilson et al. (2023), advancing bias mitigation techniques, aligning datasets with evolving LLM technologies, exploring scalable dataset maintenance solutions, and broadening the spectrum of LLM applications. Integral to this journey is the development of comprehensive ethical guidelines and the fostering of collaborative research, both of which are pivotal for the responsible evolution of LLM technology.